OpenStudy (anonymous):

Complex Functions: Cauchy's Integral Formula compute: integral of (dz/(z-3)^4 ) dz where c: /z-2/=4

OpenStudy (amistre64):

i cant make sense of how you notated it

OpenStudy (amistre64):

c is a contour im assuming

OpenStudy (amistre64):

is the dz iterated twice by accident?

OpenStudy (anonymous):

ill attach a file of the question more clearly

OpenStudy (amistre64):

that is clearer, thanx :)

OpenStudy (anonymous):

no thank you for taking the time to look at my problem

OpenStudy (amistre64):

im reading up on what i might have forgotten, c: defines a circle such that |z-2|=4

OpenStudy (anonymous):

yes thats right, and i know theres another part to it to find the answer using : 2Pi/n! multiplied by f^n (z0)

OpenStudy (amistre64):

this might take me a few ...

OpenStudy (anonymous):

its okay

OpenStudy (amistre64):

i think that n! is related to this right?

OpenStudy (anonymous):

yes i know that the n in the problem has to do with the exponent 4 . its something like this :

OpenStudy (anonymous):

yes thats right thats part of solving it

OpenStudy (amistre64):

 is this what |z-2| = 3 represents? I spose I never really did get a grasp on this ....

OpenStudy (amistre64):

= 4 that is

OpenStudy (amistre64):

\[\frac{1}{(z-3)^4}\] the 3 there tells us that theres is a point "3" inside the circle?

OpenStudy (anonymous):

i think that part is to make sure that the z was in the circle , being g that -2 is a point in the circle. I just get quite confused with this section in complex/

OpenStudy (amistre64):

\[f^{(n)}(\alpha)=\frac{1}{2\pi\ i\ }\int_{C}\frac{f(z)}{(z-\alpha)^{n+1}}\] is what im looking at

OpenStudy (amistre64):

http://ndp.jct.ac.il/tutorials/complex/node36.html

OpenStudy (amistre64):

as far as I can tell; f(z) = 1 according to that ...

OpenStudy (amistre64):

does:\[\int_C\to\int_{0}^{2\pi}\] ??

OpenStudy (anonymous):

yes i tried working on it myself a bit just just now.(z0)= -3, f(z)=1, f(z0) = -3 not quite sure if this is correct.

OpenStudy (amistre64):

\[\frac{1}{(z-3)^4}=\frac{a}{(z-3)}+\frac{b}{(z-3)^2}+\frac{c}{(z-3)^3}+\frac{d}{(z-3)^4}\] \[1=a(z-3)^3+b(z-3)^2+c(z-3)+d\] d=1 \[1=a(z-3)^3+b(z-3)^2+c(z-3)+1\] \[0=a(z^3-9z^2+27z-27)+b(z^2-6z+9)+c(z-3)\] \[0=(a)z^3+(-9a+b)z^2+(27a-6b+c)z -(27a+9b-3c)\] a=0; b=0; \[0=(c)z -(27a+9b-3c)\] c=0 sooo, this seems to be alot of working just to get :) \[2\pi\ i\left(\frac{d^3}{dz^3}(1)_{z=3}\right) \]

OpenStudy (anonymous):

i must say it is a lot to grasp. But I thank you anyways for spending a long time on it!! lol

OpenStudy (amistre64):

well, i do need the practice ;)

OpenStudy (anonymous):

i need lots of it. lol